# The Exact Variance of the Average Treatment Effect Estimator in Cluster RCT

### v1

@@ 3.1 The estimator
Our goal in this section is to calculate the exact variance of $\widehat{\tau}$ and then propose a feasible and consistent estimator of that variance for inference purposes.
To simplify notation and account for heterogeneity in cluster sizes, define the average cluster size
@@
This form makes explicit the two distinct sources of randomness in the estimator: between-cluster randomness, which is due to sampling and treatment assignment of clusters; within-cluster randomness, which is due to the second-stage sampling of units used to construct $\widehat{\widetilde{Y}}_{c}$.
@@ 4 Asymptotic Analysis
Conditional on the cluster $c$ being sampled, $n_{c}>0$.
@@ 5 Estimators of Exact Variance
$\mathrm{Var}(\widehat{\tau})\leq\mathrm{Var}(\widehat{\tau})^{\text{consv}}$. If the fourth moment of the potential outcomes is bounded, and $\left(\frac{C}{S_{d}}-1\right)\sigma^{2}\left(\widetilde{Y}(d)\right)\nrightarrow 0$, $d=0,1$, then $\widehat{\mathrm{Var}}(\widehat{\tau})^{\text{consv}}$ is an unbiased and consistent estimator of $\mathrm{Var}(\widehat{\tau})^{\text{consv}}$.
@@ 6 Simulation Results
For each unit $i$ in cluster $c$, we generate two independent covariates $X_{ci}=(X_{1ci},X_{2ci})\in(U[0,1])^{2}$, and let $\gamma=(1,1)$. We consider four distinct data generating processes that vary in how heterogeneity and clustering enter the model. They are presented in Table 1. All noise terms are Guassian with variance $25$.
@@
DGP 1 is the benchmark case, corresponding to the constant treatment effect, with all units sharing the same treatment effect $\tau=50$. DGP 2 corresponds to the case of heterogeneous treatment effects across individuals without clustering. Treatment effects vary randomly across individuals, $\tau_{i}\sim\mathcal{N}(50,100)$, but there are no cluster-level shocks. This mimics a standard i.i.d. setting. DGP 3 depicts cluster-level heterogeneity. Each cluster $c$ has its own random intercept $\alpha_{c}\sim\mathcal{N}(5,25)$, and cluster-specific treatment effect $\tau_{c}\sim\mathcal{N}(20,100)$. Units in the same cluster share the same treatment effect, but treatment effects differ across clusters. DGP 4 reflects both clustered and within-cluster heterogeneity. Clusters differ in their mean effect $\tau_{c}\sim\mathcal{N}(20,100)$, and within each cluster, individuals differ further via an idiosyncratic spread $\sigma_{c}^{2}\sim U[0,4]$, and thus have effect $\tau_{ci}\sim\mathcal{N}(\tau_{c},\sigma_{c}^{2})$. This is the most complex and realistic design, reflecting cases such as school-level interventions with student-specific idiosyncrasies.
@@
When treatment effects vary across clusters (DGP 3) or both within and across clusters (DGP 4), differences emerge: the Liang–Zeger SEs are consistently larger, leading to over-coverage ($\approx 98\%$) and lower power ($\approx 82–96\%$); the upper-bound SEs remain close to the empirical dispersion, yielding coverage near the nominal $95–97\%$ and higher power (up to +7 percentage points). This behavior aligns with theory: the LZ estimator is known to be conservative under design-based randomization, while our bound-based estimator incorporates the attainable negative covariance term, reducing unnecessary conservatism. Comparing R1 (proportional sampling) and R2 (fixed size): Proportional sampling typically yields smaller standard errors and more stable estimates because sampling intensity scales with cluster size. Fixed-size sampling oversamples smaller clusters relative to their size, which slightly inflates the total variance. The performance patterns of both variance estimators are similar across regimes, but proportional sampling tends to produce marginally more precise inference.
@@ 7 Application
Women should go to college 0.070 0.035 [0.001,0.140] 0.039 [-0.006,0.147] 8.9%
@@ 8 Conclusion
This paper develops a design-based framework for analyzing cluster randomized experiments (cluster randomized controlled trials, CRCTs) when both cluster-level assignment and two-stage sampling—of clusters and, within them, individual units—are present. We derive the exact finite-population variance of the Horvitz–Thompson estimator for the average treatment effect (ATE) and establish its asymptotic normality under mild assumptions on outcome moments, sampling probabilities, and cluster-size heterogeneity.
@@
Our simulations and empirical application, using the school-based CRCT of dhar2022reshaping, show that the proposed upper-bound variance estimator delivers tighter confidence intervals than the conventional Liang–Zeger cluster-robust estimator, while maintaining correct coverage. The improvement is most pronounced when treatment effects are heterogeneous and clustered, where standard clustered standard errors tend to be conservative. In such cases, our estimator captures the attainable negative covariance components that standard approaches ignore, leading to more efficient inference and higher power.
In summary, the paper contributes (i) exact design-based variance formulas for cluster-level treatment with two-stage sampling, (ii) sharp and conservative variance estimators that are consistent under the design, and (iii) asymptotic results linking finite-population theory to practical inference. These tools enable researchers to conduct design-faithful and efficient uncertainty quantification in cluster randomized experiments.

### v2

# The Exact Variance of the Average Treatment Effect Estimator in Cluster Randomized Controlled Trials

@@ 3.1 The estimator
Remark 1.
The difference-in-means estimator is
$\widehat{\tau}^{dm}=\frac{1}{n_{1}}\sum_{c=1}^{C}\sum_{i=1}^{N_{c}}R_{c}R_{i|c}D_{c}Y_{ci}-\frac{1}{n_{0}}\sum_{c=1}^{C}\sum_{i=1}^{N_{c}}R_{c}R_{i|c}(1-D_{c})Y_{ci},$
where
$n_{1}=\sum_{c,i}R_{c}R_{i|c}D_{c},\qquad n_{0}=\sum_{c,i}R_{c}R_{i|c}(1-D_{c})$
are the realized numbers of sampled treated and sampled control units. Thus, both denominators $n_{1}$ and $n_{0}$ are random.
By contrast, the HT estimator divides by the fixed population size $N$ and places all randomness in the sampling and treatment indicators $(R_{c},R_{i|c},D_{c})$ inside the numerator. The randomness in the realized treatment and control sample sizes is already accounted for through these indicators. Because each observed outcome is weighted by inverse inclusion and treatment probabilities, the HT estimator reconstructs the population total directly, without introducing additional randomness through a random denominator.
Our goal in this section is to calculate the exact variance of $\widehat{\tau}$ and then propose a feasible and consistent estimator of that variance for inference purposes. To simplify notation and account for heterogeneity in cluster sizes, define the average cluster size
@@
This form makes explicit the two distinct sources of randomness in the estimator: between-cluster randomness, which is due to sampling and treatment assignment of clusters, and within-cluster randomness, which is due to the second-stage sampling of units used to construct $\widehat{\widetilde{Y}}_{c}$.
@@ 4 Asymptotic Analysis
The second-stage sampling probability is not 0: $n_{c}>0$, $\forall c$.
@@ 5 Estimators of Exact Variance
$C\mathrm{Var}(\widehat{\tau})\leq C\mathrm{Var}(\widehat{\tau})^{\text{consv}}$. If the fourth moment of the potential outcomes is bounded, and $\left(\frac{C}{S_{d}}-1\right)\sigma^{2}\left(\widetilde{Y}(d)\right)\nrightarrow 0$, $d=0,1$, then $C\widehat{\mathrm{Var}}(\widehat{\tau})^{\text{consv}}$ is an unbiased and consistent estimator of $C\mathrm{Var}(\widehat{\tau})^{\text{consv}}$.
@@ 6 Simulation Results
For each unit $i$ in cluster $c$, we generate two independent covariates $X_{ci}=(X_{1ci},X_{2ci})\in(U[0,1])^{2}$ and set $\gamma=(1,1)$. We consider four data generating processes (DGPs) that differ in the extent and source of treatment effect heterogeneity and clustering. All idiosyncratic noise terms are Gaussian with variance $25$. The DGPs are summarized in Table 1.
@@
DGP 1 is a benchmark with a constant treatment effect $\tau=50$. DGP 2 introduces individual-level heterogeneity without clustering, with $\tau_{i}\sim\mathcal{N}(50,100)$. DGP 3 incorporates cluster-level heterogeneity through a cluster-specific intercept $\alpha_{c}\sim\mathcal{N}(5,25)$ and a cluster-level treatment effect $\tau_{c}\sim\mathcal{N}(20,100)$; units within the same cluster share the same treatment effect, but effects vary across clusters. DGP 4 allows for both between- and within-cluster heterogeneity: cluster means vary as $\tau_{c}\sim\mathcal{N}(20,100)$, and individual treatment effects vary within clusters as $\tau_{ci}\sim\mathcal{N}(\tau_{c},\sigma_{c}^{2})$ with $\sigma_{c}^{2}\sim U[0,4]$.
The four DGPs span a range of environments with increasing heterogeneity and clustering. DGP 1 is a benchmark with homogeneous treatment effects, serving as a baseline in which all variance estimators coincide. DGP 2 introduces individual-level treatment effect heterogeneity without clustering, corresponding to a standard i.i.d. setting. DGP 3 allows for cluster-level heterogeneity through cluster-specific treatment effects, capturing settings such as school- or village-level interventions where outcomes are correlated within clusters. DGP 4 further incorporates within-cluster heterogeneity by allowing individual treatment effects to vary around a cluster-specific mean. This final design is the most flexible and empirically realistic, combining both between- and within-cluster variation in treatment effects.
@@
When treatment effects vary across clusters (DGP 3) or both within and across clusters (DGP 4), differences emerge: the Liang–Zeger SEs are consistently larger, leading to over-coverage ($\approx 98\%$) and lower power ($\approx 82–96\%$); the upper-bound SEs remain close to the empirical dispersion, yielding coverage near the nominal $95–97\%$ and higher power (up to +7 percentage points). This behavior aligns with theory: the LZ estimator is known to be conservative under design-based randomization, while our bound-based estimator incorporates the attainable negative covariance term, reducing unnecessary conservatism.
When comparing R1 (proportional sampling) and R2 (fixed size), we can see that proportional sampling typically yields smaller standard errors and more stable estimates because sampling intensity scales with cluster size, and fixed-size sampling oversamples smaller clusters relative to their size, which slightly inflates the total variance. The performance patterns of both variance estimators are similar across regimes, but proportional sampling tends to produce marginally more precise inference.
@@ 7 Application
Women should attend college 0.070 0.035 [0.001,0.140] 0.039 [-0.006,0.147] 8.9%
@@ 8 Conclusion
This paper develops a design-based framework for analyzing cluster randomized controlled trials when both cluster-level assignment and two-stage sampling (sampling of clusters and sampling individual units within clusters) are present. We derive the exact finite-population variance of the Horvitz–Thompson estimator for the average treatment effect and establish its asymptotic normality under mild assumptions on outcome moments, sampling probabilities, and cluster-size heterogeneity.
@@
Our simulations and empirical application, using the school-based cluster randomized controlled trials of dhar2022reshaping, show that the proposed upper-bound variance estimator delivers tighter confidence intervals than the conventional Liang–Zeger cluster-robust estimator, while maintaining correct coverage. The improvement is most pronounced when treatment effects are heterogeneous and clustered, where standard clustered standard errors tend to be conservative. In such cases, our estimator captures the attainable negative covariance components that standard approaches ignore, leading to more efficient inference and higher power.
In summary, the paper contributes exact design-based variance formulas for cluster-level treatment with two-stage sampling, sharp and conservative variance estimators that are consistent under the design, and asymptotic results linking finite-population theory to practical inference. These tools enable researchers to conduct design-faithful and efficient uncertainty quantification in cluster randomized controlled trials.